can you please help me solve this equation 2/3 / 4/3 / 7/8 this how it comes out in the book someone else helped me but i still don't understand how they solved it. he said something about needing parenthesis. like this (2/3)/(4/3)/(7/8)but I'm still clueless.
Thank you

Answer by Alan3354(69443)

(2/3)/(4/3)/(7/8)
----------------
This could be [(2/3)/(4/3)]/(7/8) or (2/3)/[(4/3)/(7/8)]
----------------
[(2/3)/(4/3)]/(7/8) = (1/2)/(7/8) = 1/2*8/7 = 8/14 = 4/7
----------------
(2/3)/[(4/3)/(7/8)] = (2/3)/(4/3*8/7) = (2/3)/(32/21) = 2/3*21/32 = 21/32
You have to clarify what the problem using parentheses.
Plus, it's not an equation, there's no equal sign in (2/3)/(4/3)/(7/8)
Equal, equate, equation.
--------------------
Things written in the book are not all on the same line, so that's not how it is in the book. Saying that's how it is in the book doesn't eliminate the ambiguity.